OpenStudy (anonymous):

simplify 9x^(2/3)/3x^(1/6)

hartnn (hartnn):

Use this \(\huge \frac{x^a}{x^b}=x^{a-b}\)

hartnn (hartnn):

and as u know 9/3 will be 3.

OpenStudy (lgbasallote):

unless....

hartnn (hartnn):

unless the question actually is, (9x^(2/3)/3)x^(1/6)

OpenStudy (anonymous):

\[9x ^{2/3}\div 3x ^{1/6}\]

OpenStudy (anonymous):

its 3x^(1/2)...as power of x are subtracted 2/3 - 1/6.

hartnn (hartnn):

so what u got as \(\huge \frac{x^{2/3}}{x^1/6}=x^{2/3-1/6}=?\)

OpenStudy (anonymous):

so I subtract as I normally would for fractions?

hartnn (hartnn):

yup, what will be 2/3-1/6 ?

OpenStudy (anonymous):

4/6-1/6=3/6=1/2. I See now.

OpenStudy (anonymous):

thanks to everybody that posted a response.

hartnn (hartnn):

welcome :)